A drug used to treat asthma is 55% bioavailable as 5-mg tablets to be taken once daily. If a patient is switched to the inhalant form of the drug, which is 87% bioavailable, how many metered 500-ug sprays should the patient administer every 12 hours to receive an equivalent drug dose? The answer is 3.16 or 3 sprays. Please show how this question is solved.
Answer #1

Bioavailability of the drug in tablet form=55% of 5 mg

= 55*5000/100

=2750micro grams

Bioavailability of drug in inhaler form=500*87/100

=435micrograms

So if the person consumes 5 mg tablet he will get 2750micrograms of drug in bioavailable form

For equivalent amount of drug number of metered sprays required per day (24 hrs) =2750/435=6.32

So every 12 hours number of inhalers =3.16